### v1

# On existence of neutral graph

Abstract: Graph is considered neutral if its assortativity coefficient $r$ is equal to zero. In this paper, we address an outstanding conjecture, i.e., whether is there a neutral graph on $n$ vertices. First, we show that for $n\geq 7$, there is at least one neutral tree, which suggests that we find a representative of any order neutral graph. Additionally, we obtain that given $n\geq 13$, there exist at least one neutral non-tree graph.
Keywords: Neutral graph, Assortativity coefficient.
1 Introduction
Let $G=(V,E)$ represent a simple connected graph, then, notations $V$ and $E$, respectively, indicates vertex set and edge set. We use $G$ as shorthand of $G=(V,E)$ when it is clear from the context. The number of vertices (a.k.a., order) of graph $G$ is $n(:=|V|)$, and the total number of edges in graph $G$ is $m(:=|E|)$. Assorataivity coefficient $r$ of graph $G$ is defined as follows [2]
$r=\frac{m^{-1}\sum\limits_{e_{uv}\in E}d_{u}d_{v}-\left[m^{-1}\sum\limits_{e_{uv}\in E}\frac{1}{2}(d_{u}+d_{v})\right]^{2}}{m^{-1}\sum\limits_{e_{uv}\in E}\frac{1}{2}(d^{2}_{u}+d^{2}_{v})-\left[m^{-1}\sum\limits_{e_{uv}\in E}\frac{1}{2}(d_{u}+d_{v})\right]^{2}},$ (1)
in which $d_{u}$ denotes degree of vertex $u$ and $e_{uv}$ represents an edge whose endpoints are vertices $u$ and $v$, separately. Graph $G$ is considered assortative if $r>0$, disassortative if $r<0$, and neutral otherwise. Obviously, it is easy to determine whether a given graph is neutral or not. However, to the best of our knowledge, there have not been published work to answer the following conjecture [3]-[6]
Is there neutral graph of arbitrary order?
In this work, we provide a positive answer to the aforementioned issue.
2 Definition and main results
This section comprises two subsections as follows.
2.1 Definition
First of all, we introduce some notations and terminologies used throughout in this work. Note that graph under consideration is simple and connected.
Definition 1[7] A graph $G$ is considered $k$-regular if and only if each vertex $u$ has degree $d_{u}=k$. For simplicity, we denote by $k$-$G$ $k$-regular graph. $k$-$G$ is a cycle if $k$ is equal to $2$.
Definition 2[7] The $s$-th subdivision graph $G_{s}=(V_{s},E_{s})$ is obtained from graph $G$ by inserting $s$ vertices into each edge.
Definition 3 The single-edge division graph $G^{\prime}=(V^{\prime},E^{\prime})$ is obtained from graph $G$ by only inserting one vertex into an arbitrarily chose edge.
Definition 4 The triangle-operation is to (i) first add an isolated vertex for each edge of graph $G$, and (2) connect each isolated vertex to two endvertices of the associated edge using two edges. The resulting graph is referred to as $G^{\bigtriangleup}=(V^{\bigtriangleup},E^{\bigtriangleup})$ for brevity.
Definition 5 The leaf-connecting-operation is to (1) connect $d_{v}$ leaf vertices to each vertex $v$ in graph $G$, and (2) connect a pair of leaves using an edge, then perform the same connection between another two leaves until no leaf is left. The resulting graph is referred to as $G^{\dagger}=(V^{\dagger},E^{\dagger})$ for presentation. Degree one vertex is defined as a leaf.
2.2 Main results
In the sequel, let us show main results composed of two subsections. The former shows existence of neutral tree, and the existence of neutral non-tree graph is displayed in the later.
2.2.1 Neutral tree
Here we prove the existence of neutral tree.
Lemma 1 Graph $G$ is either a cycle of any order or a tree with a unique largest degree $\Delta=3$ vertex, if the following expression holds
$\sum_{e_{uv}\in E}(d_{u}+d_{v})=4m.$ (2)
From which it follows that the desired tree has order $n\geq 4$.
Proof It is easy to prove Lemma 1 using some elementary computations. ∎
Lemma 2 Given graph $G$,
$2m\sum\limits_{e_{uv}\in E}(d^{2}_{u}+d^{2}_{v})-\left[\sum\limits_{e_{uv}\in E}(d_{u}+d_{v})\right]^{2}\geq 0.$ (3)
Equality holds iff $G$ is regular.
Proof By means of Cauchy-Schwartz inequality, we prove Lemma 2. ∎
Based on Lemma 2, it is clear to see that an irregular graph $G$ is neutral iff the numerator of Eq. (1) is equal to zero.
Lemma 3 In $s$-th subdivision graph $G_{s}=(V_{s},E_{s})$,
$4|E_{s}|\sum\limits_{e_{ij}\in E_{s}}d_{i}d_{j}-\left[\sum\limits_{e_{ij}\in E_{s}}(d_{i}+d_{j})\right]^{2}=-\left[\sum_{e_{uv}\in E}(d_{u}+d_{v}-4)\right]^{2}.$ (4)
Proof After some simple algebra, we complete Lemma 3. ∎
Theorem 1 Given any tree $T$ subjected to Eq.(2), the corresponding $s$-th subdivision tree $T_{s}$ is neutral.
Proof Using Lemmas 1 and 3, we verify Theorem 1. ∎
According to Theorem 1, together with Lemma 1, any tree of order 2n12n-1 turns out to be neutral by setting $s=1$. In addition, using simple calculations, we reach the following corollary.
Corollary 1 Any tree $T^{\prime}_{s}$ is neutral, and its single-edge division tree $(T^{\prime}_{s})^{\prime}$ is neutral as well.
Now, based on Theorem 1 and Corollary 1, it is easy to prove that there is at least one neutral tree of order $n+3$. Additionally, we easily verify, by enumeration, that any graph is not neutral if its order is no more than $6$. Armed with the results above, we firmly state that the next theorem holds true.
Theorem 2 There exists at least one tree of order $n+3$ such that the associated assortativity coefficient $r$ is equivalent to zero.
By far, we succeed in finding a representative of any order neutral graph. It should be mentioned that the representative has tree structure. Next, we are concerned with other representatives having non-tree structure.
2.2.2 Neutral non-tree graph
Below elaborates on the existence of neutral non-tree graph.
Lemma 4 Given an arbitrary neutral tree $T$ on $n_{1}(\geq 7)$ vertices, then the graph $T^{\dagger}$ is neutral.
Proof After some simple algebra, we complete Lemma 4. ∎
Clearly, neutral graph $T^{\dagger}$ is not a tree. It should be mentioned that graph $T^{\dagger}$ has $3n_{1}-2$ vertices.
Lemma 5 Given an arbitrary cycle $G$ on $n_{2}(\geq 3)$ vertices, then the graph $G^{\dagger}$ is neutral.
Proof As before, with some simple algebra, we complete Lemma 5. ∎
Similarly, neutral graph $G^{\dagger}$ is not a tree either. At the same time, graph $G^{\dagger}$ has $3n_{2}$ vertices.
Lemma 6 Given an arbitrary neutral tree $T$ on $n_{1}(\geq 7)$ vertices, then the graph $T^{\bigtriangleup}$ is neutral.
Proof Analogously, with some simple algebra, we complete Lemma 6. ∎
It is clear to see that non-tree graph $T^{\bigtriangleup}$ has $2n_{1}-1$ vertices.
Until now, we obtain that there is a neutral non-tree graph where the total number of vertices $n$ obeys
$n\mod 6=0,1,3,4,5,\qquad n\geq\max\left\{\min\{3n_{1}-2\},\min\{3n_{2}\},\min\{2n_{1}-1\}\right\}=19$ (5)
Next, we will prove case $n\mod 6=2$. To this end, we need introduce a class of graphs. Without loss of generality, we use $G^{\oplus}$ to represent these graphs. Graph $G^{\oplus}$ is obtained based on two small graphs. Concretely speaking, given a neutral graph $G(1)=(V(1),E(1))$ and a $k$-regular graph $G(2)=(V(2),E(2))$, then we require that (i) $\beta$ incomplete edges [^2] An edge is considered incomplete if it just connects a vertex. be stemmed from each vertex of graph $G(2)$ and (ii) for each edge of graph $G(1)$, $(\alpha-1)$ incomplete edges be stemmed from every endvertex. In the meantime, we assume that $2(\alpha-1)|E(1)|=\beta|V(2)|$ holds true. Next, one merges an incomplete edge from graph $G(1)$ with another incomplete edge in graph $G(2)$ into an edge. Performing such manipulation until no incomplete edge is left. The resultant graph is $G^{\oplus}$.
Lemma 7 Graph $G^{\oplus}$ is neutral if the following expression holds
$\frac{|E(2)|}{|E(1)|}=(\alpha-1)^{2}=\left(\frac{k}{\beta}\right)^{2}=\left(\frac{|V(2)|}{|V(1)|}\right)^{2}.$ (6)
Proof By definition in (1), we have
$\sum\limits_{e_{ij}\in E^{\oplus}}d_{i}d_{j}$ $=\sum\limits_{e_{uv}\in E(1)}\alpha^{2}d_{u}d_{v}+\sum\limits_{u\in V(1)}\alpha d_{u}(k+\beta)(\alpha-1)d_{u}+(k+\beta)^{2}|E(2)|$ (7)
$=\sum\limits_{e_{uv}\in E(1)}\alpha^{2}d_{u}d_{v}+\alpha(\alpha-1)(k+\beta)\sum\limits_{u\in V(1)}d_{u}^{2}+(k+\beta)^{2}|E(2)|.$
$\sum\limits_{e_{ij}\in E^{\oplus}}d_{i}+d_{j}$ $=\sum\limits_{e_{uv}\in E(1)}(\alpha d_{u}+\alpha d_{v})+\sum\limits_{u\in V(1)}[\alpha d_{u}+(k+\beta)](\alpha-1)d_{u}$ (8)
$\quad+2(k+\beta)|E(2)|$
$=\alpha^{2}\sum\limits_{e_{uv}\in E(1)}(d_{u}+d_{v})+2(k+\beta)[(\alpha-1)|E(1)|+|E(2)|].$
Next, it is easy to check
$4|E^{\oplus}|\left(\sum\limits_{e_{ij}\in E^{\oplus}}d_{i}d_{j}\right)-\left(\sum\limits_{e_{ij}\in E^{\oplus}}d_{i}+d_{j}\right)^{2}$ (9)
$=\alpha^{4}\left[4|E(1)|\left(\sum\limits_{e_{uv}\in E(1)}d_{u}d_{v}\right)-\left(\sum\limits_{e_{uv}\in E(1)}d_{u}+d_{v}\right)^{2}\right]$
$\equiv 0,$
in which $|E^{\oplus}|=(2\alpha-1)|E(1)|+|E(2)|$. This is complete. ∎
From the results above, we obtain
$|V^{\oplus}|=|V(1)|+|V(2)|=\alpha|V(1)|.$
After that, if we suppose that parameter $\alpha$ is equal to $2$, and neutral graph $G(1)$ is graph $T^{\dagger}$, the resulting graph $G^{\oplus}$ satisfies
$|V^{\oplus}|=2(3n_{1}-2)\mod 6=2.$
To sum up, we come to the following theorem.
Theorem 3 Given $n\geq 38$, there exists at least one neutral non-tree graph of order $n$.
As a byproduct, we have the following assertion about graph $G^{\ominus}$. The $G^{\ominus}$ is obtained from an arbitrary neutral graph $G$ and its copy $G^{\circ}$ in a similar manner as used in the construction of graph $G^{\oplus}$. Concretely speaking, one connects each vertex $u$ in graph $G$ to neighboring vertices of its peer $u^{\circ}$ in copy $G^{\circ}$.
Assertion 1 Graph $G^{\ominus}$ shares the same assortativity coefficient with graph $G$.
Proof By definition in (1), we have
[…]
$\sum\limits_{e_{ij}\in E^{\ominus}}d_{i}+d_{j}$ $=4\sum\limits_{e_{uv}\in E}(2d_{u}+2d_{v})=8\sum\limits_{e_{uv}\in E}(d_{u}+d_{v}),$ (11)
[…]
Next, it is easy to check
$4|E^{\ominus}|$ $\left(\sum\limits_{e_{ij}\in E^{\ominus}}d_{i}d_{j}\right)-\left(\sum\limits_{e_{ij}\in E^{\ominus}}d_{i}+d_{j}\right)^{2}$ (13)
[…]
and
$2|E^{\ominus}|$ $\left[\sum\limits_{e_{ij}\in E^{\ominus}}(d^{2}_{i}+d^{2}_{j})\right]-\left(\sum\limits_{e_{ij}\in E^{\ominus}}d_{i}+d_{j}\right)^{2}$ (14)
[…]
in which we have used quality $|E^{\ominus}|=4|E|$. By far, it is easy to prove $r^{\ominus}\equiv r$. This is complete. ∎
Obviously, if we suppose that graph $G$ is graph $T^{\dagger}$, the resulting graph $G^{\ominus}$ also satisfies
$|V^{\ominus}|=2(3n_{1}-2)\mod 6=2.$
In addition, if an arbitrary tree $T$ on $n_{1}(\geq 7)$ vertices stated in Theorem 2 is selected as graph to create graph $G^{\ominus}$, then the total number of vertices of the end graph $G^{\ominus}$ is equal $2n_{1}$. Together with Lemma 6, we reach the last theorem in this work as follows.
Theorem 4 Given $n\geq 13$, there exists at least one neutral non-tree graph of order $n$.
Acknowledgments
The research was supported in part by the National Natural Science Foundation of China No. 62403381, the Key Research and Development Plan of Shaanxi Province No. 2024GX-YBXM-021, and the Fundamental Research Funds for the Central Universities No. G2025KY06223